• Summer 2014 TEM course

    January 23, 2014. We were unable to offer more than 8 seats for the spring TEM course. As a result, some students were unable to register. To fulfill the need we are considering a summer offering if there is enough interest. Please reply if you would be interested in the …

    JEOL JEM-2010 TEM
  • New TEM online now

    January 14, 2014. The new 2010 TEM is being used now by experienced users. Operation of the 2010 is very similar to the 4000FX (old TEM but there are still some differences I need to demonstrate. Please contact me when you need to use it. We will be installing the …

    JEOL JEM-2010 TEM
  • TEM EDS Detector

    January 13, 2014. The EDS detector for the TEM will be installed Wednesday. The TEM will likely be down until Friday morning.

    JEOL JEM-2010 TEM
  • TEM Status

    December 3, 2013. The TEM is operational and the digital camera is attached and working. I need to calibrate the magnification and camera length next Monday morning. I should be able to start limited training later next week. The vibration isolation equipment will be installed during the week of Dec …

    JEOL JEM-2010 TEM
  • TEM update

    November 17, 2013. The service engineer will return this week to complete installation of the new(er) JEOL 2010 TEM. Key parts of this effort include stabilizing the accelerating voltage and burning in the new cathode. Ideally we want to move the high speed digital camera from the 4000 TEM to …

    JEOL JEM-2010 TEM

Ferromagnetic particles in FESEM

April 22, 2014. We noticed that the astigmatism could not be adequately corrected even with the stigmator adjustment at the end stop. I opened the specimen chamber and found ferromagnetic particles attached to the objective lens by the magnetic field. This is not good. It makes trouble for everyone. We …

Owen vacation

March 21, 2014. I will be on vacation beginning Wed, April 2 through Mon, April 7. The labs will be open and …


FESEM stage control

March 4, 2014. I reset the stage control and the motion is correct now as well as the drag-to-center option in the …


Labs closed

February 21, 2014. The university closed at 2:00. We are closing the labs after the current users finish their work. All later …


FESEM open now

February 20, 2014. FESEM open now and can be used today. Remember that the chamber camera view has changed.


FESEM backscatter detector

February 20, 2014. Yesterday we installed an accessory BSE detector on the FESEM. Unfortunately, there is a problem with the electronics and …


FESEM is down

February 3, 2014. A sample has been dropped in the chamber and blocks anyone else from using the SEM until it is …